How would I find the IRR of the investment given the following information:

​Pondside Properties is evaluating a real estate investment of Oceanflow Estates. Management plans to buy the property today and sell it 5 years from today. The initial cost of the property is $7 million and the expected sale price is $13 million (answer as percentage to 2 decimal places).
Answer #1

Let irr be x%
At irr,present value of inflows=present value of outflows.

7=13/1.0x^5

1.0x^5=(13/7)

1.0x=(13/7)^(1/5)

x=1.1318-1

Hence x=irr=13.18%(Approx).